OpenStudy (anonymous):

@dumbcow @jim_thompson5910 @ganeshie8 Which shows three ordered pairs that are all solutions of the equation y = x + 4? A. (1, 4), (2, 6), and (3, 7) B. (1, 4), (2, 8), and (3, 12) C. (3, 7), (5, 9), and (7, 10) D. (3, 7), (5, 9), and (7, 11)

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

Plug in the x and y's from your solution and if you get y = x then that's your solution.

OpenStudy (xguardians):

Has to be C or D since if you input (1,4) it incorrect. And only D works, all the inputs go in. D is answer:) Just input x and y into the equation.

OpenStudy (anonymous):

@xGuardians Don't give the answers, only lead them to the right direction.

jimthompson5910 (jim_thompson5910):

Checking (1,4) y = x + 4 4 = 1 + 4 4 = 5 ... false so (1,4) is not a solution. You do the same for the others

OpenStudy (anonymous):

Which three ordered pairs complete the table to give solutions of the equation y = x – 7? A. (–3, –10), (0, –7), and (3, 4) B. (–3, –10), (0, –7), and (3, –4) C. (–3, –4), (0, –7), and (3, –4) D. (–3, 4), (0, –7), and (3, –4) http://static.k12.com/calms_media/media/1417500_1418000/1417734/1/b311ca753cda78973c0eeef8f3548665f70934fb/FGA_130919_161117.jpg

OpenStudy (anonymous):

@xGuardians plz help me

OpenStudy (anonymous):

Do the same thing

OpenStudy (anonymous):

From your following solutions (a,b,c,d) plus in the (x,y) coordinates into the function and see if the left side equals the right side.

OpenStudy (anonymous):

Jim did a example for you

OpenStudy (anonymous):

Y=x-7

OpenStudy (anonymous):

Yeah that's your function, now use the coordinates from the options, and test each one to see which is correct.

OpenStudy (anonymous):

For example, let's do A, I'll do the first coordinate, then you have to do the rest. (-3, -10) y = x-7 -10=-3-7 -10 = -10 So this is right, now do it with the rest of them.

OpenStudy (anonymous):

@xGuardians help plz

OpenStudy (xguardians):

Answer is D.

OpenStudy (anonymous):

Well it seems like you don't care about the process @Hotchellerae21 you had everything on how to do the problem, yet you ignore it. So, I wish you the best in your future studies.